OpenStudy (anonymous):

Completely factor using the idea of GCF. 16x^3-24x^2-36x

OpenStudy (anonymous):

First look at the coefficients : 16, -24 and -36. Can you find a common factor? Next look at the variables: x^3, x^2 and x can you find a common factor?

OpenStudy (anonymous):

I got 4 but I'm not sure how to find the variable...

OpenStudy (anonymous):

Alright well, look at it this way: x^3=x*x*x x^2=x*x and x=x

OpenStudy (anonymous):

well don't they all have only 1 x in common?

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

So your GCF must be 4x

OpenStudy (anonymous):

Oh ok i get it now! thank you!

OpenStudy (anonymous):

No prob :)